The linear charge density of an infinite line charge is 0.3 10 .m C − −16  Assuming 

that the electric potential at a perpendicular distance of 5.0 m from the wire is zero, 

calculate the potential at the perpendicular distance of 6.0 m.


1
Expert's answer
2021-06-21T16:02:50-0400

The potential at distance r from infinite charged wire is


"V=\frac{\lambda}{2\pi \epsilon_0}\ln r"

"V=\frac{0.3\cdot10^{-3}}{2\pi (8.85\cdot10^{-12})}\ln 6=9.7\ MV"